OpenStudy (anonymous):

algebra 1b solve by factoring 3x squared + 7x-6=0

OpenStudy (anonymous):

(3x - 2) (x + 2) = 0

OpenStudy (anonymous):

i mean (3x-2) ( x + 3)

OpenStudy (anonymous):

\[3x ^{2}+7x-6\]

OpenStudy (vijay):

x=2/3 and -3

OpenStudy (anonymous):

how do you get that. i really need to understand

OpenStudy (anonymous):

trial and error you can also use quadratic equation

OpenStudy (vijay):

3x^2 + 7x-6=0 3x^2 + 9x-2x-6=0 3x(x+3)-2(x+3)=0 (x+3)(3x-2)=0

OpenStudy (anonymous):

\[step 1. (x+9)(x-2) step 2??\]

OpenStudy (vijay):

(x+3)(3x-2)=0 (x+3) = 0 and 3x-2 = 0 x = -3 and x = 2/3

OpenStudy (anonymous):

how did you get (x+3)(3x-2) ????

OpenStudy (vijay):

from this line 3x(x+3)-2(x+3)=0 if u factorise (x+3) , u may get

OpenStudy (anonymous):

correction: how did you get \[3x ^{2}+7x-6=0\]

OpenStudy (anonymous):

sorry im confusing myself .

OpenStudy (anonymous):

my teacher makes us do a big x. so what multiplies to -12 and adds to 7. i got 9 and -2 so (x+9) (x-2) how do i go on from there

OpenStudy (vijay):

can u type your question again with mathematical symbols ?

OpenStudy (anonymous):

solve by factoring \[3x ^{2}+7x-6=0\] step one. FACTOR. multiply -6 and 3 to get -18. now find what multiplies to -18 make sure that they also add to get 7. 9 and -2 MULTIPLY together and make -18 and ADD to make 7. so you make the equation (x+9)(x-2)=0 but you still have to do more for the rest of the problem. so im confused on the rest

OpenStudy (vijay):

substitute your -9 and 2 for 7

OpenStudy (vijay):

sorry 9 and -2

OpenStudy (anonymous):

so it would be \[3x ^{2}+(x+9)(x-2)-6=0 ?\] that makes no sense

OpenStudy (vijay):

no, \[3x^2 +9x - 2x -6=0\]

OpenStudy (anonymous):

so what do i do to solve that? add like terms?

OpenStudy (vijay):

then,. factorise 3x from from 1st two terms ans factorise 2 from last two terms

OpenStudy (anonymous):

3x(x+3) -2(x+3)

OpenStudy (vijay):

yes, then factorise (x+3)

OpenStudy (anonymous):

wait so how to i solve the rest of the problem

OpenStudy (vijay):

then, each of them equal to zero

OpenStudy (anonymous):

3x(x+3)=0 -2(x+3)=0

OpenStudy (vijay):

no, when u factorise by ( x-3), u will get (x-3)(3x-2) = 0

OpenStudy (vijay):

After that make, x-3 = 0 and 3x- 2 = 0

OpenStudy (anonymous):

howd you get (x-3) and (3x-2)

OpenStudy (vijay):

from 3x(x+3) -2(x+3) = 0, factorise (x+3)

OpenStudy (vijay):

take out x+3 , then 3x-2 is other balance

OpenStudy (anonymous):

how do you get 3x-2

OpenStudy (vijay):

from 3x(x+3) -2(x+3), take out x-3, from 1st term 3x and 2nd term -2 is balance

OpenStudy (anonymous):

because (x+3)(x+3) cancel out. leaving ONE (x+3) and whats left is 3x-2?

OpenStudy (vijay):

3x(x+3) -2(x+3) if u see front of (x+3),there is 3x and -2.

OpenStudy (anonymous):

3x-2=0 x-2=0/3 x=3/1+2/1 x=6 x+3=0 x= -3 x= 3 x= 6

OpenStudy (vijay):

nope, 3x-2 = 0 and x+3=0 3x=2 x=-3 x=3/2

OpenStudy (anonymous):

not x=2/3 ???

OpenStudy (vijay):

x = 2/3 is correct, if u substitute in to the function, it will satisfy.

OpenStudy (anonymous):

ok so x= 2/3 x= -3

OpenStudy (vijay):

yes, that is the final answer, 100% correct

OpenStudy (anonymous):

okay( thank you SOOO much!